A piece of lumber of fixed length ?? is broken into two sub pieces at position ?? ~ ??(0,??). Let ?? denote the length of the shorter sub piece.
a. Find the cumulative distribution function of ??. Hint: First express ?? as a piecewise function in terms of ??.
b. Find the probability distribution function of ??. Identify the name of the distribution that ?? follows. Also write down the parameter value(s) of this distribution.
Answer #1

Since X\sim U\left ( 0,l \right ) , then Y is defined as Y=\left\{\begin{matrix} X, & 0<X\leqslant \frac{l}{2}\\ l-X, & \frac{l}{2}<X<l \end{matrix}\right.

a) Now

For 0<X\leqslant \frac{l}{2}

P\left ( Y<y \right )=P\left ( X<y \right )\\ P\left ( Y<y \right )=\frac{y}{l},0<y<\frac{l}{2}

For \frac{l}{2}<X<l

P\left ( Y<y \right )=P\left ( l-X<y \right )\\ P\left ( Y<y \right )=P\left ( X>l-y \right )\\ P\left ( Y<y \right )=1-P\left ( X\leqslant l-y \right )\\ P\left ( Y<y \right )=1-\frac{l-y}{l}\\ P\left ( Y<y \right )=\frac{y}{l},0<y<\frac{l}{2}

Thus, the CDF of Y is {\color{Blue} F_Y\left ( y \right )=\frac{2y}{l},0<y<\frac{l}{2}}

b) The pdf is

f_Y\left ( y \right )=\frac{\mathrm{d} }{\mathrm{d} y} F_Y\left ( y \right )\\ {\color{Blue} f_Y\left ( y \right )=\frac{2}{l},0<y<\frac{l}{2}}

The distribution of Y is uniform in \left ( 0,\frac{l}{2} \right ) . That is {\color{Blue} Y\sim \left ( 0,\frac{l}{2} \right )}